OpenStudy (anonymous):

the square root of X-6 - the square root of X = 3

OpenStudy (anonymous):

im thinking we can probably square both sides of the equation

OpenStudy (dumbcow):

yep square both sides, you will have to FOIL left side [sqrt(x-6) -sqrt(x)]^2 = 3^2

OpenStudy (dumbcow):

actually it may be easier to move the sqrt(x) to right side first before squaring both sides sqrt(x-6)^2 = [3+sqrt(x)]^2

OpenStudy (anonymous):

ok, so i got x-6= 9+x. is this when i foil?

OpenStudy (dumbcow):

No FOIL the [3+sqrt(x)]^2 it does NOT equal 9+x (3+sqrt(x))(3+sqrt(x)) = 9+3sqrt(x) +3sqrt(x) +x

OpenStudy (anonymous):

does this need up being a no solution problem?

OpenStudy (dumbcow):

anyway, the answer is No solution sqrt(x-6) < sqrt(x) sqrt(x-6) - sqrt(x) < 0 a number less than zero can never equal 3

OpenStudy (anonymous):

Thank you for the help!